Mature salmon swim upstream, returning to spawn at their birthplace. During the arduous trip they leap vertically upward over waterfalls as high as 2.81 m. With what minimum speed (in m/s) must a salmon launch itself into the air to clear a 2.81-m waterfall?
Answer #1

Using 3rd equation of motion,

V^2 = u^2 + 2gh

0 = u^2 - (2 x 9.8 x 2.81)

Velocity, u = 7.42 m/s

Comment in case any doubt please rate my answer....